# Defending against adversarial attacks using mixture of experts

### v1

@@ 5 Experiments
Table 2: Accuracy and robustness of our proposed defense system vs. other models in terms of standard accuracy (SA) and robust accuracy (RA).
Model Metric SA RA - FGSM RA - PGD
0.01 0.02 0.03 0.04 0.05 0.06 0.07 0.08 0.09 10 20 30 40 50
VGG-19 93.18 52.08 50.92 48.52 41.53 29.22 20.22 16.06 13.70 12.00 42.66 41.91 41.72 41.80 41.80
ResNet-18 92.35 52.32 48.19 40.97 30.28 21.38 15.61 12.72 11.61 11.00 37.72 37.28 37.10 37.02 36.95
ResNet-50 95.29 52.68 51.54 46.01 35.04 25.46 19.26 16.25 13.87 12.14 46.20 45.68 45.53 45.44 45.45
DenseNet-169 95.10 55.32 54.57 48.63 36.38 25.21 18.57 15.00 12.88 11.90 47.55 46.99 46.81 46.72 46.81
GoogLeNet 95.01 53.83 51.53 46.96 33.59 22.70 16.61 13.83 11.69 10.91 39.60 38.59 38.53 38.48 38.49
Inception-V3 94.78 58.00 56.77 52.05 41.99 29.99 21.89 16.53 13.61 11.97 47.21 45.51 45.01 44.68 44.80
Xception 93.48 51.41 50.20 47.69 37.72 26.35 18.45 14.21 12.32 10.96 41.36 40.58 40.44 40.30 40.30
MobileNet-V2 94.20 55.05 53.79 49.31 36.67 24.70 18.31 15.38 13.39 12.19 45.77 45.20 45.14 45.17 45.08
SoE 81.69 70.23 66.22 60.27 52.65 44.62 38.26 32.96 28.74 25.74 62.27 62.34 62.31 62.31 62.31
ADVMoE 88.27 79.39 62.09 40.88 23.90 16.53 15.47 15.98 16.07 16.40 47.69 44.95 44.36 44.11 44.01
DWF (Ours) 91.08 82.83 79.66 74.76 69.01 63.60 58.40 52.97 47.34 41.24 65.51 64.7 64.52 64.42 64.37
@@
All experiments were conducted on a server equipped with NVIDIA H100 GPU, offering the adequate computational power for MoE training. We implemented all the code in PyTorch with torch 2.7.1 and torchvision 0.22.0, and CUDA 12.8.
Dataset
To conduct our experiments, we use CIFAR-10 [krizhevsky2009learning] dataset, which is an RGB-colored dataset, including 10 labels.
Models
Figure 5: The backbone architecture of DWF network.
For every expert we use ResNet-18 [he2016deep] as the backbone architecture, since other relevant defenses use the same backbone. Figure 5 shows the backbone details. To compare our results with plain networks we used pretrained networks that are used in [liu2017pytorchcifar, juraev2022cifar10models]. These models include VGG-19 [simonyan2014very], ResNet-50 [he2016deep], DenseNet-169 [huang2017densely], GoogLeNet, Inception-V3 [szegedy2016rethinking], Xception [chollet2017xception], and MobileNet-V2 [sandler2018mobilenetv2]. The attained results are shown in Table 2.
@@
Table 3: Attacks settings.
@@
To evaluate our model’s resilience against white box attacks, we employ FGSM with the perturbation sizes $\epsilon\in\{0.01,~0.02,~\dots,~0.09\}$, and PGD with the step size $\alpha=\frac{2}{255}$, perturbation bound $\epsilon=\frac{8}{255}$, and iterations numbers $i\in\{10,~20,~30,~40,~50\}$. These setting were used during both training and test phases, highlighting that both threat model and defense system have white-box knowledge about each other. Table 3 shows an overview of the attack settings.
@@
where $c\in\{R,G,B\}$ stands for each color channel, $x^{\prime}_{c}$ is the normalized version of $x_{c}$, $\mu_{c}$ is the empirical mean over channel $c$, $\sigma_{c}$ is the empirical deviation over channel $c$, and $i,j$ are row index and column index, respectively. To normalize the data we defined $\mu=(0.4914,~0.4822,~0.4465)$ and $\sigma=(0.2023,~0.1994,~0.2010)$.
@@
Additionally, to further improve our model’s stability, we employed a cosine annealing learning rate scheduler [loshchilov2016sgdr] to decay the learning rate gradually over all iterations in all training sessions. Table 4 shows an overview of the hyper parameters used for training our proposed approach.
ADVMoE used TRADES [zhang2019theoretically] as the adversarial training objective by adopting PGD with attack strength of $\epsilon=\frac{8}{255}$ during two iterations. Additionally, it was trained during 200 epochs with learning rate of 0.1, momentum of 0.9, weight decay of $5e^{-4}$, and batch size of 128. SoE was adversarially trained during 200 epochs with learning rate of 0.1 and batch size of 128. All the experiments were done using the same backbone (ResNet-18) for DWF, ADVMoE, and SoE to ensure a fair comparison.
@@ Results
Table 2 illustrates the accuracy and robustness of our proposed model in comparison with other models. Figures 6 and 7 present a visual summary of the results reported in Table 2 and provide clearer comparison between our proposed model and other competitors in terms of robust accuracy under different settings for FGSM and PGD attacks. The attained results by our model are averaged over multiple runs, ensuring a fair comparison. Figures 8 and 9 show the statistics over these runs. The results over these runs vary slightly due to the randomness during training sessions. This randomness occurs due to four reasons: i) training data shuffling, ii) random weight initialization of the models before training, iii) random FGSM perturbation size of adversarial sample generation for each batch in every epoch, and iv) random PGD iteration of adversarial sample generation for each batch in every epoch.
Moreover, Figures 8 and 9 show that our proposed approach suffers from accuracy degradation, when faces stronger attacks, which is very common among all other competitors. We can see that our proposed model outperforms all other competitors in terms of robustness under all attacks. Additionally, DWF consistently outperforms existing defenses such as ADVMoE and SoE in terms of clean accuracy. Notably, it maintains a clean accuracy above $90\%$ ($91.08\pm 0.39$), highlighting its effectiveness not only against adversarial attacks but also in non-adversarial environments. Another insight illustrates that as we increase perturbation size or iterations in FGSM and PGD, the robustness of all other models drops, which shows the lack of adaptability of the models against attacks with more extreme specifications.
Figure 8: Box plots showing the robustness of our proposed approach against PGD w.r.t. different iteration numbers over multiple runs.
Figure 9: Box plots showing the robustness of our proposed approach against FGSM w.r.t. different perturbation sizes over multiple runs.
@@ 6 Conclusion and Future Works
In this paper, we proposed a novel defense system combining the ensemble architecture of MoE with adversarial training to robustify our model against PGD and FGSM attacks. Our method outperforms other competitors including undefended networks and state-of-the-art defense systems. Moreover, our proposed method, unlike the other defense systems, does not extremely sacrifice clean accuracy, underscoring its success in both adversarial and non-adversarial environments.
To reduce training time, efficient adversarial training method such as FFAT can be used. However, efficient adversarial training reduces the robustness of the model. We also trained and tested our model against two popular attacks, FGSM and PGD. To further expand our method one can train and evaluate the model on a wider range of adversarial attacks.

### v2

@@ 5 Experiments
Table 2: Accuracy and robustness of our proposed defense system vs. other models in terms of standard accuracy (SA) and robust accuracy (RA) (FGSM) - DenseNet-169 (DN169).
Model Metric SA RA - FGSM
0.01 0.02 0.03 0.04 0.05 0.06 0.07 0.08 0.09
CIFAR-10
VGG-19 93.18 52.08 50.92 48.52 41.53 29.22 20.22 16.06 13.70 12.00
ResNet-18 92.35 52.32 48.19 40.97 30.28 21.38 15.61 12.72 11.61 11.00
ResNet-18 (FGSM-AT) 89.22 78.81 67.50 48.37 32.71 23.52 16.48 12.2 9.54 9.26
ResNet-18 (PGD-AT) 87.99 78.83 64.29 41.46 28.34 20.8 15.29 11.57 9.8 9.78
ResNet-50 95.29 52.68 51.54 46.01 35.04 25.46 19.26 16.25 13.87 12.14
DN169 95.10 55.32 54.57 48.63 36.38 25.21 18.57 15.00 12.88 11.90
GoogLeNet 95.01 53.83 51.53 46.96 33.59 22.70 16.61 13.83 11.69 10.91
Inception-V3 94.78 58.00 56.77 52.05 41.99 29.99 21.89 16.53 13.61 11.97
Xception 93.48 51.41 50.20 47.69 37.72 26.35 18.45 14.21 12.32 10.96
MobileNet-V2 94.20 55.05 53.79 49.31 36.67 24.70 18.31 15.38 13.39 12.19
SoE 81.69 70.23 66.22 60.27 52.65 44.62 38.26 32.96 28.74 25.74
ADVMoE 88.27 79.39 62.09 40.88 23.90 16.53 15.47 15.98 16.07 16.40
[…]
SVHN
ResNet-18 95.95 70.98 67.85 63.03 54.87 46.20 35.59 25.96 18.29 13.77
ResNet-34 96.26 70.54 67.2 63.12 57.53 48.74 38.74 28.38 19.89 14.6
ResNet-50 96.32 69.02 66.75 63.03 57.98 50.81 41.39 31.33 22.6 16.62
DenseNet-121 95.53 70.36 66.87 61.52 56.2 48.94 40.74 32.36 25.09 19.97
ADVMoE 88.88 79.22 62.87 41.69 24.71 15.86 14.05 14.05 13.47 13.42
[…]
Table 3: Accuracy and robustness of our proposed defense system vs. other models in terms of PGD robust accuracy.
Model Metric RA - PGD
10 20 30 40 50 100 (MS) 100 (A)
CIFAR-10
VGG-19 42.66 41.91 41.72 41.80 41.80 41.52 41.7
ResNet-18 37.72 37.28 37.10 37.02 36.95 37.07 37
ResNet-18 (FGSM-AT) 58.65 58.21 58.17 58.14 58.13 58 58.11
ResNet-18 (PGD-AT) 58.74 58.35 58.31 58.26 58.26 58.05 58.26
ResNet-50 46.20 45.68 45.53 45.44 45.45 44.86 45.39
DenseNet-169 47.55 46.99 46.81 46.72 46.81 45.72 46.73
GoogLeNet 39.60 38.59 38.53 38.48 38.49 38.34 38.39
Inception-V3 47.21 45.51 45.01 44.68 44.80 43.66 44.55
Xception 41.36 40.58 40.44 40.30 40.30 40.01 40.31
MobileNet-V2 45.77 45.20 45.14 45.17 45.08 45.03 45.12
SoE 62.27 62.34 62.31 62.31 62.31 62.03 62.27
ADVMoE 47.69 44.95 44.36 44.11 44.01 42.47 43.49
[…]
SVHN
ResNet-18 49.61 48.86 48.62 48.56 48.52 48.29 48.46
ResNet-34 49.04 48.04 47.83 47.71 47.66 47.37 47.54
ResNet-50 48.74 47.88 47.7 47.6 47.54 47.25 47.45
DenseNet-121 45.94 44.54 44.29 44.18 44.14 43.82 44.05
ADVMoE 47.22 44.22 43.31 43 42.68 8.13 8.33
[…]
@@
All experiments were conducted on a server equipped with two NVIDIA H100 GPUs each with 80 GB of VRAM, and 512 GB of RAM, offering the adequate computational power for MoE training. While we used two NVIDIA H100 GPUs to accelerate experimentation, the proposed DWF architecture is composed of lightweight ResNet-18 experts and is modular. Training can be performed on more modest GPU environments by reducing batch size and/or using gradient accumulation; additional ablations on scaling with fewer experts are left for future work. Importantly, inference is substantially cheaper than training and can be deployed on commodity GPU hardware. We implemented all the code in PyTorch with torch 2.7.1 and torchvision 0.22.0, and CUDA 12.8.
DWF introduces additional computational overhead compared to single-network baselines due to the joint training and inference of multiple experts. On our hardware, full training requires several days (approximately five days of wall-clock time, corresponding to roughly 240 GPU-hours), whereas single baselines complete substantially faster. ADVMoE and SoE exhibit intermediate training costs between these extremes. During inference, the gating module is lightweight and the dominant cost comes from evaluating the experts, resulting in approximately linear scaling with the number of experts. Although DWF is computationally heavier than plain networks, we find the overhead acceptable in practice given the considerable robustness gains. This highlights a typical robustness-efficiency trade-off.
Datasets
To conduct our experiments, we use CIFAR-10 [28] and street view house number (SVHN) [38] datasets, which are RGB-colored datasets, each include 10 labels.
@@
Moreover, to further assess the resilience of our proposed approach, we conduct further experiments experiments under white-box multi-step PGD and adaptive PGD attacks.
Table 4: Attacks settings - expectation over transformation (EoT).
@@
Multi-Step PGD $\epsilon=\frac{8}{255}$, $\alpha=\frac{2}{255}$, $i=100$, $\text{Restart}=3$ white-box black-box
Adaptive PGD $\epsilon=\frac{8}{255}$, $\alpha=\frac{2}{255}$, $i=100$, $\text{EoT Samples}=1$ white-box black-box
@@
Benign baselines
To compare our results with plain networks we used networks pretrained on CIFAR-10 from [34, 24] and networks pretrained on SVHN from [14]. These models include VGG-19 [51], ResNet-34, ResNet-50 [19], DenseNet-121, DenseNet-169 [21], GoogLeNet, Inception-V3 [53], Xception [12], and MobileNet-V2 [48].
Plain-network baselines are initialized from standard publicly available pretrained checkpoints. Although their original training recipes may differ slightly, all models are evaluated under an identical test-time protocol, including the same preprocessing, normalization, and attack settings, ensuring controlled robustness comparisons. These checkpoints achieve strong clean accuracy, indicating that they are competitive baselines rather than undertrained variants. The substantial robustness gap observed between our method and these baselines suggests that minor differences in the original training procedures are unlikely to affect the main conclusions. The attained results are shown in Table 3.
Adversarially trained baselines
To evaluate whether the observed resilience stems from the adversarial training schedule, the ensemble structure, or a combination of both, we conducted adversarial training experiments with two separate ResNet-18 models using different adversarial training methods: FGSM and PGD.
Backbone
For every expert we use ResNet-18 [19] as the backbone architecture, since other relevant defenses use the same backbone. Figure 5 shows the backbone details.
Figure 5: The backbone architecture of DWF network.
@@
We adopt a full white-box threat model where the adversary has complete knowledge of the entire DWF architecture, including all experts, the gating network, and routing weights. Gradients are computed through the full mixture-of-experts pipeline, allowing attacks to jointly optimize perturbations with respect to both expert predictions and gating behavior. Therefore, the attacker explicitly targets the complete MoE routing mechanism rather than individual experts only. To further approximate worst-case adversarial conditions, we additionally evaluate stronger multi-step and adaptive attacks that account for input-dependent routing, ensuring that the robustness assessment is not affected by gradient masking or routing-specific obfuscation.
To evaluate our model’s resilience against white box attacks, we employ FGSM with the perturbation sizes $\epsilon\in\{0.01,~0.02,~\dots,~0.09\}$, and PGD with the step size $\alpha=\frac{2}{255}$, perturbation bound $\epsilon=\frac{8}{255}$, and iterations numbers $i\in\{10,~20,~30,~40,~50\}$. These setting were used during both training and test phases, highlighting that both threat model and defense system have white-box knowledge about each other. Moreover, to further evaluate our proposed defense’s performance against unseen attacks (black-box defense knowledge), we used multi-step PGD-100 with 3 restarts and adaptive PGD-100 with expectation over transformation (EoT) of one. Table 4 shows an overview of the attacks settings.
@@
where $c\in\{R,G,B\}$ stands for each color channel, $x^{\prime}_{c}$ is the normalized version of $x_{c}$, $\mu_{c}$ is the empirical mean over channel $c$, $\sigma_{c}$ is the empirical deviation over channel $c$, and $i,j$ are row index and column index, respectively. To normalize CIFAR-10 data we defined $\mu=(0.4377,~0.4438,~0.4728)$ and $\sigma=(0.1980,~0.2010,~0.1970)$. Moreover, to normalize SVHN data we defined $\mu=(0.4914,~0.4822,~0.4465)$ and $\sigma=(0.2023,~0.1994,~0.2010)$.
@@
Additionally, to further improve our model’s stability, we employed a cosine annealing learning rate scheduler [35] to decay the learning rate gradually over all iterations in all training sessions. Table 5 shows an overview of the hyper parameters used for training our proposed approach.
ADVMoE used TRADES [58] as the adversarial training objective by adopting PGD with attack strength of $\epsilon=\frac{8}{255}$ during two iterations. Additionally, it was trained during 200 epochs with learning rate of 0.1, momentum of 0.9, weight decay of $5e^{-4}$, and batch size of 128. SoE was adversarially trained during 200 epochs with learning rate of 0.1 and batch size of 128. The first ResNet-18 model was adversarially trained using the FGSM with a perturbation magnitude of $\epsilon=0.03$, where half of each training batch consisted of FGSM-generated adversarial examples. The second ResNet-18 model was trained using PGD, with the adversarial perturbation set to $\epsilon=\frac{8}{255}$, step size $\alpha=\frac{2}{255}$, and 10 iteration, where half of each training batch consisted of PGD-generated adversarial examples.
To ensure fair comparisons, all methods use the same backbone architecture (ResNet-18), optimizer, learning rate schedule, and training protocol. Some baselines (e.g., ADVMoE and SoE) also employ multiple experts, although the number of experts varies across methods as part of their architectural design. A larger number of experts increases computational cost, analogous to increasing network width or depth, and should therefore be interpreted as a capacity trade-off rather than an advantage in training settings. Our comparisons keep the optimization procedure identical while allowing each method to retain its original architecture.
@@ Results
Table 3 illustrates the accuracy and robustness of our proposed model in comparison with other models. Figures 6 and 7 present a visual summary of the results reported in Table 3 and provide clearer comparison between our proposed model and other competitors in terms of robust accuracy under different settings for FGSM and PGD attacks. The attained results by our model are averaged over five runs, ensuring a fair comparison.
To isolate architectural effects from training-related factors, all compared defended methods, including ADVMoE and SoE, are trained and evaluated under the same experimental pipeline with identical preprocessing, backbone configuration, and optimization protocol. Robust accuracy is computed using the same threat models for every method. Consequently, performance differences primarily reflect differences in architectural design rather than inconsistencies in training or evaluation procedures.
The results over these runs vary slightly due to the randomness during training sessions. This randomness occurs due to four reasons: i) training data shuffling, ii) random weight initialization of the models before training, iii) random FGSM perturbation size of adversarial sample generation for each batch in every epoch, and iv) random PGD iteration of adversarial sample generation for each batch in every epoch.
Moreover, our proposed approach suffers from robust accuracy degradation, when faces stronger attacks, which is very common among all other competitors. We can see that our proposed model outperforms all other competitors in terms of robustness under all attacks. Additionally, DWF consistently outperforms existing defenses such as ADVMoE, SoE, and adversarially trained networks in terms of clean accuracy. Notably, it maintains a clean accuracy above 90% across both datasets, highlighting its effectiveness not only against adversarial attacks but also in non-adversarial environments. Compared to plain (undefended) networks, DWF exhibits at most a modest reduction in clean accuracy, which is expected since DWF incorporates adversarial training and jointly optimizes robustness-oriented objectives through the gating mechanism and expert updates. This optimization slightly shifts capacity toward worst-case robustness, whereas plain models are optimized purely for benign performance. Importantly, the clean accuracy drop in DWF is noticeably smaller than in standard adversarially trained baselines, indicating that the mixture-of-experts design helps preserve discriminative capacity on benign inputs while still providing substantial robustness gains.
Moreover, the results demonstrate that our method could preserve its performance under adaptive white-box scenario, suggesting that the resilience does not stem from gradient masking/obfuscation. Another insight illustrates that as we increase perturbation size or iterations in FGSM and PGD, the robustness of all other models drops, which shows the lack of adaptability of the models against attacks with more extreme specifications. As a result, it is useful to consider how the gating mechanism may behave under stronger perturbations. In a soft-gated mixture-of-experts architecture, routing can distribute weights across multiple experts when predictions become uncertain, which may help avoid complete reliance on any single expert and lead to gradual performance degradation as attack strength increases. We focus on overall robustness trends in this work.
The robustness gains of DWF can be interpreted through several complementary mechanisms. First, multiple experts introduce functional diversity, which reduces correlated failure modes and makes adversarial transfer across experts more difficult. Second, soft gating enables input-dependent specialization, allowing each expert to focus on a subset of the data manifold and improving local decision boundaries. Third, joint optimization with adversarial training encourages smoother and more stable gradients across the mixture. Together, these effects contribute to improved robustness without a large sacrifice in clean accuracy.
@@ 6 Conclusion and Future Works
In this study, we evaluated DWF’s robustness primarily under white-box attacks, which represent the strongest and most conservative threat model since the adversary has full access to the model parameters and gradients. Robustness under white-box optimization is widely regarded as a lower bound on performance under weaker settings, such as transfer-based or query-limited black-box attacks, which typically yield higher robust accuracy. Therefore, defenses that remain effective in the white-box scenario are expected to generalize well to these weaker threat models. A systematic evaluation of DWF under transfer and black-box attacks is an interesting direction for future work.
DWF integrates several complementary mechanisms, including soft gating, fractional expert activation, joint expert-gating optimization, and randomized adversarial training. Each component addresses a distinct limitation of conventional defenses. Soft gating enables input-dependent expert specialization, fractional activation reduces overfitting and improves diversity, joint updating ensures coordinated optimization between experts and routing, and randomized adversarial training improves robustness across varying perturbation strengths. Together, these components contribute synergistically to both clean and robust accuracy. Conducting a full combinatorial ablation over all variants would require retraining numerous MoE configurations with substantial computational cost; therefore, we leave a comprehensive ablation study for future work.
While this work focuses primarily on improving robustness, training efficiency is also an important practical consideration. Fast adversarial training variants (e.g., FFAT or single-step approximations) may reduce computational cost at the expense of some robustness. A systematic study quantifying the trade-off between training time and robust accuracy for DWF under different efficient training strategies would require multiple full training runs and extensive hyperparameter tuning. We leave this investigation for future work. Moreover, We trained and tested our model against two popular attacks, FGSM and PGD. To further expand our method one can train and evaluate the model on a wider range of adversarial attacks.
Although the proposed framework is evaluated on image classification tasks, the soft-gated mixture-of-experts design is largely architecture-agnostic and can, in principle, be integrated with other modalities such as text, speech, or time-series models. The core idea of training specialized experts and adaptively routing inputs through a gating network does not depend on convolutional backbones. However, extending the method to other domains may require modality-specific adversarial threat models and training strategies, and could introduce additional computational costs. Exploring these extensions is an interesting direction for future work.